# Geometry of horospheres in Kobayashi hyperbolic domains

### v1

@@ 1. Introduction and statement of the results
$H^{s}_{o}(x,R)$ :=$:=\Big{\{}z\in\Omega\,:\,\limsup_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big{)}<(1/2)\log R\Big{\}}$
$H^{b}_{o}(x,R)$ :=$:=\Big{\{}z\in\Omega\,:\,\liminf_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big{)}<(1/2)\log R\Big{\}}.$
@@
In the case $\Omega=\mathbb{D}$, the open unit disc in the complex plane centered at 00, using the formula for $\mathsf{k}_{\mathbb{D}}$, a straightforward computation shows that when $o=0$, $x\in\partial\mathbb{D}$, for each $z\in\mathbb{D}$, one has
@@
Geometrically, $H(x,R)$ is the disc with center $x/{(1+R)}$ and radius $R/(R+1)$, internally tangent to $\partial\mathbb{D}$ at $x$. In particular, closures of both the small and big horospheres of arbitrary radius with center $x$ and pole at 00 intersect the boundary of the unit disc only at the point $x$. An analogous result holds true for the Euclidean unit ball $\mathbb{B}^{n}$ of $\mathbb{C}^{n}$, $n\geq 2$. On the other hand, for the $n$-dimensional polydisc ${\mathbb{D}}^{n}$, $n\geq 2$, things are quite different. In this case, taking $o=0\in{\mathbb{D}}^{n}$, Abate showed that $\overline{H^{s}_{0}(x,R)}\cap\partial{\mathbb{D}}^{n}=\{x\}$ if and only if $x$ is on the Shilov boundary $(\partial\mathbb{D})^{n}$ of ${\mathbb{D}}^{n}$. On the other hand, $\overline{H^{b}_{0}(x,R)}\cap\partial{\mathbb{D}}^{n}$ always contains $\{x\}$ properly (see Proposition 2.4.12 and Corollary 2.4.13 in [2]).
@@
As an upshot of the above result and the existence of limit in (1.2), for the three classes of domains considered by Arosio et al., property (1.1) in the problem ()(*)( ∗ ) holds true. However, whether (1.1) holds true for a general model-Gromov-hyperbolic domain is not known. In this article, we answer this question affirmatively. Before we state this result, a notation: since we would also like to consider unbounded domains, the natural formulation of model-Gromov-hyperbolic domains is via the end-point compactification $\overline{\Omega}^{End}$. For bounded domains $\overline{\Omega}^{End}=\overline{\Omega}$; see [14] for more details about end-point compactification.
@@
Owing to the aforementioned result of Bharali–Zimmer and Proposition 1.2, we know that $\overline{H^{b}_{o}(x,R)}\cap\partial\Omega=\{x\}$ for every $x\in\partial\Omega$, for a model-Gromov-hyperbolic domain. Gromov hyperbolicity plays a crucial role in showing that $x\in\overline{H^{s}_{o}(x,R)}$. More specifically, it is the property of uniform separation of geodesic rays starting at a fixed point in the domain and landing at the same point in the Gromov boundary that is at the heart of our proof that $\overline{H^{s}_{o}(x,R)}\cap\partial\Omega=\{x\}$.
There is no complete characterization of a model-Gromov-hyperbolic domain. Except the three considered by Arosio et al., a few others are also listed in [11].
A natural question arises whether there are non-Gromov-hyperbolic domains having property (1.1) in the problem ()(*)( ∗ ). We provide two results in this direction. Our first result deals with domains that are locally model-Gromov-hyperbolic domain; see Section 4 for the definition.
@@
Let $\Omega\subset\mathbb{C}^{d}$ be an complete Kobayashi hyperbolic domain that is hyperbolically embedded in $\mathbb{C}^{d}$. Suppose $\Omega$ is locally model-Gromov-hyperbolic domain. Then for any $x\in\Omega$, $o\in\Omega$ and $R>0$, we have
@@
The domains in Theorem 1.3, Theorem 1.4 and Theorem 1.5 are all visible domains. It is natural to ask if the visibility property itself implies the property (1.1) in problem ()(*)( ∗ ). This is not true, e.g., the domain $\Omega:=\mathbb{D}\setminus\{[0,1)\}$ is a visibility domain but for any $x\in(0,1]$, there exists an $R_{0}$, depending on $x$, such that for any $R<R_{0}$, we have
@@
The Euclidean unit ball $\mathbb{B}^{d}\subset\mathbb{C}^{d}$, $d\geq 1$, is metrically-regular as one explicitly knows the formulae for the Kobayashi distance. On the other hand, $\mathbb{D}^{d}$, $d\geq 2$, is not metrically-regular. The three types of domains considered by Arosio et al. are metrically-regular. We are now ready to present our main result.
@@
Note that any bounded domain that satisfies condition (1.1) in problem ()(*)( ∗ ) satisfies the condition $(a)$ and $(b)$ imposed on the domain $\Omega$ in the statement of the above theorem. In particular, taking $D=\mathbb{D}$ and $\Omega$ to be a bounded simply-connected domain in $\mathbb{C}$ that satisfies (1.1) we recover Theorem 1.1. Theorem 1.7 also implies that bounded simply-connected domains that are metrically-regular are the Jordan domains.
@@ 3. General results about horospheres
Lemma 3.3 also helps in understanding why for the bidisc ${\mathbb{D}}^{2}$, given $x=(x_{1},x_{2})\in\partial{{\mathbb{D}}^{2}}$, and $R>0$, $H^{b}_{o}(x,R)$, with pole 00, contains $x$ properly — as observed by Abate by direct computation. To see how, assuming $x_{1}\in\partial\mathbb{D}$, we consider a geodesic ray of ${\mathbb{D}}^{2}$ of the form $\gamma:=(\sigma,f\circ\sigma)$, where $\sigma$ is the radial geodesic ray of the unit disc starting at $0\in\mathbb{D}$ and landing at $x_{1}$ and $f$ is a holomorphic self-map of $\mathbb{D}$ such that the radial cluster set of $f$ at $x_{1}$ contains $x_{2}$ properly. It is a non-trivial fact that such an $f$ exist. Clearly then $\mathcal{C}(\gamma)$ contains $(x_{1},x_{2})$ properly and then by the above observation $H^{b}_{o}(x,R)$ contains $(x_{1},x_{2})$ properly.
@@
Result 3.4 (paraphrasing Zimmer [24, Lemma 3.2]).
@@
Let $\gamma_{n}:[0,L_{n}]\longrightarrow\Omega$ be a geodesic segment joining $o$ and $w_{n}$ and let $\sigma_{n}:[0,M_{n}]\longrightarrow\Omega$ be a geodesic segment joining $z_{T}$ and $w_{n}$. Appealing to Ascoli–Arzelá Theorem, and passing to subsequences, without loss of any generality, we can assume that $\gamma_{n}$, $\sigma_{n}$ converges to geodesic rays $\gamma$ and $\sigma$, respectively, locally uniformly on $[0,\infty)$. Since $\Omega$ is a visibility domain, by [10, Lemma 3.1], both $\gamma$ and $\sigma$ land at the point $x$. By our assumption, $x$ represents a unique point in the Gromov boundary $\partial_{G}\Omega$, therefore both $\gamma$ and $\beta$ represent the same point in the Gromov boundary as represented by $x$. Since $\Omega$ is $\delta$-hyperbolic, using [12, Lemma 3.3, Chapter III.H], we have
@@ 5. Continuous extension of biholomorphisms
In this section, we present the proof of our result regarding the continuous extension of biholomorphisms. As an upshot, we shall characterize those bounded simply-connected domains in $\mathbb{C}$ that have property (1.1) in problem ()(*)( ∗ ). We begin with a lemma.
@@
Let $D$, $\Omega$ be bounded domains that are complete with respect to the Kobayashi distance. Suppose $D$ is a metrically-regular domain and $\Omega$ has the property that for every $x\in\partial\Omega$, for every $R>0$, $H^{s}_{p}(x,R)$ is nonempty with respect to some (and hence any) pole $p$. Let $\Phi:\Omega\longrightarrow D$ be a biholomorphism. Then $\Phi$ has a continuous extension from $\overline{\Omega}$ onto $\overline{D}$.
@@
Since $D$ is metrically-regular, there exists a point $o\in D$ such that $(1)$ and $(2)$ in Definition 1.6 are satisfied. Choose $p\in\Omega$ such that $\Phi(p)=o$.
Claim. Given $x\in\partial\Omega$, let $(w_{n})\subset\Omega$ be a sequence that converges to $x$ such that $\Phi(w_{n})\to\xi\in\partial D$. Then
$\Phi(H^{s}_{p}(x,R))\subset H(\xi,R)\ \ \ \forall R>0.$
To see the claim, let $z\in H^{s}_{p}(x,R)$, then by definition
$\limsup_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(z,w_{n})-\mathsf{k}_{\Omega}(o,w_{n})\big{)}\leq\limsup_{w\to x}\big{(}\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big{)}<(1/2)\log R$
where $(w_{n})$ is as in the statement of the claim. Since $\Phi$ is an isometry with respect to the Kobayashi distances, and $\Phi(w_{n})\to\xi$, the last condition implies
$\lim_{w\to\xi}\big{(}\mathsf{k}_{D}(\Phi(z),w)-\mathsf{k}_{D}(o,w)\big{)}$ $=\lim_{n\to\infty}\big{(}\mathsf{k}_{D}(\Phi(z),\Phi(w_{n}))-\mathsf{k}_{D}(\Phi(p),\Phi(w_{n})\big{)}$
$=\limsup_{n\to\infty}\big{(}\mathsf{k}_{D}(\Phi(z),\Phi(w_{n}))-\mathsf{k}_{D}(\Phi(o),\Phi(w_{n}))\big{)}$
$<(1/2)\log R.$
[…]
Now fix a point $x\in\partial\Omega$, we shall show that the limit
@@
exists. Suppose, to get a contradiction, that the above limit does not exist. This implies that — since $\Phi$ is a biholomorphism and $D$ is bounded — there exist sequences $(z_{n})$ and $(w_{n})$ in $\Omega$ that converge to $x$ and points $\xi_{1}\neq\xi_{2}\in\partial D$ such that $\Phi(z_{n})\to\xi_{1}$, $\Phi(w_{n})\to\xi_{2}$. From the above claim, we get
$\Phi(H^{s}_{o}(x,R))\subset H(\xi_{1},R)\cap H(\xi_{2},R)\ \ \ \forall R>0.$
Since $\xi_{1}$ and $\xi_{2}$ are distinct, there exists an $R_{0}>0$ such that $H(\xi_{1},R)\cap H(\xi_{2},R)=\emptyset$ for all $R<R_{0}$. This implies that $H^{s}_{o}(x,R)=\emptyset$ for all $R<R_{0}$ which is a contradiction to our hypothesis whence the result. ∎
@@
Claim. Given $\zeta\in\partial D$, let $(\zeta_{n})\subset D$ be a sequence that converges to $\zeta$ such that $\Psi(\zeta_{n})\to y$ for some $y\in\partial\Omega$. Then
@@
To see the claim, let $z\in H(\zeta,R)$, then by definition
$\lim_{w\to\zeta}\big{(}\mathsf{k}_{D}(z,w)-\mathsf{k}_{D}(o,w)\big{)}=\lim_{n\to\infty}\big{(}\mathsf{k}_{D}(z,\zeta_{n})-\mathsf{k}_{D}(o,\zeta_{n})\big{)}<(1/2)\log R,$
where $(\zeta_{n})$ is as above. Since $\Psi$ is an isometry with respect to the Kobayashi distances, and $\Psi(\zeta_{n})\to y$, the above inequality leads to
$\liminf_{w\to y}\big{(}\mathsf{k}_{\Omega}(\Psi(z),w)-\mathsf{k}_{\Omega}(p,w)\big{)}$ $\leq\liminf_{n\to\infty}\big{(}\mathsf{k}_{\Omega}(\Psi(z),\Psi(\zeta_{n}))-\mathsf{k}_{\Omega}(\Psi(o),\Psi(\zeta_{n}))\big{)}$
$=\lim_{n\to\infty}\big{(}\mathsf{k}_{D}(z,\zeta_{n})-\mathsf{k}_{D}(o,\zeta_{n})\big{)}$
$<(1/2)\log R.$
[…]
@@
Suppose, to get a contradiction, that the above is not true. This implies that — since $\Psi$ is a biholomorphism and $\Omega$ is bounded — there exist sequences $(z_{n}),(w_{n})\subset D$ converging to $\xi$ and $x_{1}\neq x_{2}\in\partial\Omega$ such that $\Psi(z_{n})\to x_{1}$ and $\Psi(w_{n})\to x_{2}$. From the above claim, we have

### v2

@@ 1. Introduction and statement of the results
$H^{s}_{o}(x,R)$ $:=\Big\{z\in\Omega\,:\,\limsup_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)<(1/2)\log R\Big\}$
$H^{b}_{o}(x,R)$ $:=\Big\{z\in\Omega\,:\,\liminf_{w\to x}\big(\mathsf{k}_{\Omega}(z,w)-\mathsf{k}_{\Omega}(o,w)\big)<(1/2)\log R\Big\}.$
@@
In the case $\Omega=\mathbb{D}$, the open unit disc in the complex plane centered at 0, using the formula for $\mathsf{k}_{\mathbb{D}}$, a straightforward computation shows that when $o=0$, $x\in\partial\mathbb{D}$, for each $z\in\mathbb{D}$, one has
@@
Geometrically, $H(x,R)$ is the disc with center x/(1+R)x/{(1+R)} and radius $R/(R+1)$, internally tangent to $\partial\mathbb{D}$ at $x$. In particular, closures of both the small and big horospheres of arbitrary radius with center $x$ and pole at 0 intersect the boundary of the unit disc only at the point $x$. An analogous result holds true for the Euclidean unit ball $\mathbb{B}^{n}$ of $\mathbb{C}^{n}$, $n\geq 2$. On the other hand, for the $n$-dimensional polydisc ${\mathbb{D}}^{n}$, $n\geq 2$, things are quite different. In this case, taking $o=0\in{\mathbb{D}}^{n}$, Abate showed that $\overline{H^{s}_{0}(x,R)}\cap\partial{\mathbb{D}}^{n}=\{x\}$ if and only if $x$ is on the Shilov boundary $(\partial\mathbb{D})^{n}$ of ${\mathbb{D}}^{n}$. On the other hand, $\overline{H^{b}_{0}(x,R)}\cap\partial{\mathbb{D}}^{n}$ always contains $\{x\}$ properly (see Proposition 2.4.12 and Corollary 2.4.13 in [2]).
@@
As an upshot of the above result and the existence of limit in (1.2), for the three classes of domains considered by Arosio et al., property (1.1) in the problem ()(*) holds true. However, whether (1.1) holds true for a general model-Gromov-hyperbolic domain is not known. In this article, we answer this question affirmatively. Before we state this result, a notation: since we would also like to consider unbounded domains, the natural formulation of model-Gromov-hyperbolic domains is via the end-point compactification $\overline{\Omega}^{End}$. For bounded domains $\overline{\Omega}^{End}=\overline{\Omega}$; see [14] for more details about end-point compactification.
@@
Owing to the aforementioned result of Bharali–Zimmer and Proposition 1.2, we know that $\overline{H^{b}_{o}(x,R)}\cap\partial\Omega=\{x\}$ for every $x\in\partial\Omega$, for a model-Gromov-hyperbolic domain. Gromov hyperbolicity plays a crucial role in showing that $x\in\overline{H^{s}_{o}(x,R)}$. More specifically, it is the property of uniform separation of geodesic rays starting at a fixed point in the domain and landing at the same point in the Gromov boundary that is at the heart of our proof that $\overline{H^{s}_{o}(x,R)}\cap\partial\Omega=\{x\}$. There is no complete characterization of a model-Gromov-hyperbolic domain. Except the three considered by Arosio et al., a few others are also listed in [11].
Remark 1.4.
We also present an alternative proof — suggested by the referee — that does not rely on visibility. This second proof draws upon the horofunction compactification, the correspondence between the horofunction and Gromov boundaries together with the uniform boundedness of horofunctions corresponding to the same Gromov boundary point. The two proofs emphasize different geometric aspects, namely — visibility versus the horofunction compactification.
A natural question arises whether there are non-Gromov-hyperbolic domains having property (1.1) in the problem ()(*). We provide two results in this direction. Our first result deals with domains that are locally model-Gromov-hyperbolic domain; see Section 4 for the definition.
@@
Let $\Omega\subset\mathbb{C}^{d}$ be an complete Kobayashi hyperbolic domain that is hyperbolically embedded in $\mathbb{C}^{d}$. Suppose $\Omega$ is locally model-Gromov-hyperbolic domain. Then for any $x\in\partial\Omega$, $o\in\Omega$ and $R>0$, we have
@@
The domains in Theorem 1.3, Theorem 1.5 and Theorem 1.6 are all visible domains. It is natural to ask if the visibility property itself implies the property (1.1) in problem ()(*). This is not true, e.g., the domain $\Omega:=\mathbb{D}\setminus\{[0,1)\}$ is a visibility domain but for any $x\in(0,1]$, there exists an $R_{0}$, depending on $x$, such that for any $R<R_{0}$, we have
@@
The Euclidean unit ball $\mathbb{B}^{d}\subset\mathbb{C}^{d}$, $d\geq 1$, is metrically-regular as one explicitly knows the formulae for the Kobayashi distance. On the other hand, $\mathbb{D}^{d}$, $d\geq 2$, is not metrically-regular. The three types of domains considered by Arosio et al. are metrically-regular.
Remark 1.8.
It is evident that when $\Omega$ is bounded, the metric regularity of $\Omega$ implies that the identity map on $\Omega$ extends as a homeomorphism from its horofunction compactification $\overline{\Omega}^{H}$ to its Euclidean closure $\overline{\Omega}$.
We are now ready to present our main result.
@@
Note that any bounded domain that satisfies condition (1.1) in problem ()(*) satisfies the condition $(a)$ and $(b)$ imposed on the domain $\Omega$ in the statement of the above theorem. In particular, taking $D=\mathbb{D}$ and $\Omega$ to be a bounded simply-connected domain in $\mathbb{C}$ that satisfies (1.1) we recover Theorem 1.1. Theorem 1.9 also implies that bounded simply-connected domains that are metrically-regular are the Jordan domains.
@@ 3. General results about horospheres
Lemma 3.3 also helps in understanding why for the bidisc ${\mathbb{D}}^{2}$, given $x=(x_{1},x_{2})\in\partial{{\mathbb{D}}^{2}}$, and $R>0$, $H^{b}_{o}(x,R)$, with pole 0, contains $x$ properly — as observed by Abate by direct computation. To see how, assuming $x_{1}\in\partial\mathbb{D}$, we consider a geodesic ray of ${\mathbb{D}}^{2}$ of the form $\gamma:=(\sigma,f\circ\sigma)$, where $\sigma$ is the radial geodesic ray of the unit disc starting at $0\in\mathbb{D}$ and landing at $x_{1}$ and $f$ is a holomorphic self-map of $\mathbb{D}$ such that the radial cluster set of $f$ at $x_{1}$ contains $x_{2}$ properly. It is a non-trivial fact that such an $f$ exist. Clearly then $\mathcal{C}(\gamma)$ contains $(x_{1},x_{2})$ properly and then by the above observation $H^{b}_{o}(x,R)$ contains $(x_{1},x_{2})$ properly.
@@
Result 3.4 (paraphrasing Zimmer [25, Lemma 3.2]).
@@
Let $\gamma_{n}:[0,L_{n}]\longrightarrow\Omega$ be a geodesic segment joining $o$ and $w_{n}$ and let $\sigma_{n}:[0,M_{n}]\longrightarrow\Omega$ be a geodesic segment joining $z_{T}$ and $w_{n}$. Appealing to Ascoli–Arzelá Theorem, and passing to subsequences, without loss of any generality, we can assume that $\gamma_{n}$, $\sigma_{n}$ converges to geodesic rays $\gamma$ and $\sigma$, respectively, locally uniformly on $0,\infty)$. Since $\Omega$ is a visibility domain, by [10, Lemma 3.1], both $\gamma$ and $\sigma$ land at the point $x$. By our assumption, $x$ represents a unique point in the Gromov boundary $\partial_{G}\Omega$, therefore both $\gamma$ and $\beta$ represent the same point in the Gromov boundary as represented by $x$. Since $\Omega$ is $\delta$-hyperbolic, using [12, Lemma 3.3, Chapter III.H], we have
@@
We now present an alternative proof of Theorem 1.3 — as alluded to in Remark 1.4. However, we first recall a few definitions, and state two results that are at the heart of this proof.
Given $(\Omega,\mathsf{k}_{\Omega})$, a complete Kobayashi hyperbolic domain, that is Gromov hyperbolic, we shall denote by $\overline{\Omega}^{H}$, $\partial\overline{\Omega}^{H}$ the horofunction compactification, and the horofunction boundary respectively. We refer the reader to [4, Section 2] for a brief introduction in our setting to these objects. Given $\xi\in\partial\overline{\Omega}^{H}$, and $p\in\Omega$, by definition, there exists a unique continuous function $h_{p,\,\xi}$, vanishing at $p$, that represents $\xi$. Given $R>0$, the set
$H_{p}(\xi,R):=\left\{x\in X\,:\,h_{p,\,\xi}(x)<\tfrac{1}{2}\log R\right\}$
is called the horoball of radius $R$ centered at $\xi$ with pole $p$. It is known fact [24, Proposition 4.6] that there exists a unique $\Phi:\overline{\Omega}^{H}\longrightarrow\overline{\Omega}^{G}$ extending the identity map continuously on $\Omega$. Moreover, we have the following results:
Result 4.1 (paraphrasing [24, Proposition 4.4]).
Let $\xi_{1}$, $\xi_{2}\in\partial\overline{\Omega}^{H}$ be such that $\Phi(\xi_{1})=\Phi(\xi_{2})$. Then, for any $p\in\Omega$,
$\big|h_{p,\,\xi_{1}}(z)-h_{p,\,\xi_{2}}(z)\big|\leq 2\delta\quad\text{for all}\,z\in\Omega,$
where $\delta\geq 0$ is such that $(\Omega,\mathsf{k}_{\Omega})$ is $\delta$-hyperbolic.
Result 4.2 (paraphrasing [4, Proposition 6.5]).
Given $\xi\in\partial\overline{\Omega}^{H}$, $p\in\Omega$, for any $R>0$,
$\overline{H_{p}(\xi,R)}^{G}\cap\partial_{G}{\Omega}=\{\Phi(\xi)\}.$
Here, the closure at the left-hand side is in the Gromov topology.
Proof.
Fix a point $o\in\Omega$, $x\in\partial\Omega$ and let $R>0$ be given. Note $x$ represents a unique point in $\partial_{G}\Omega$. We claim that there exists an $R^{\prime}>0$ such that
$H_{o}(\xi,R^{\prime})\subset H^{b}_{o}(x,R^{\prime})\subset H^{s}_{o}(x,R)\subset H_{o}(\xi,R)\quad\forall\xi\in\Phi^{-1}(x).$
To see the claim, note that it is a fact that for any $S>0$
$H^{b}_{o}(x,S)=\bigcup_{\xi\in\Phi^{-1}(x)}H_{o}(\xi,S),\quad H^{s}_{o}(x,S)=\bigcap_{\xi\in\Phi^{-1}(x)}H_{o}(\xi,S).$
Therefore, we only need to show the second inclusion in the claim above. Let $R^{\prime}>0$ be such that $(1/2)\log{R^{\prime}}+2\delta\leq{(1/2)}\log{R}$. Let $z\in H^{b}_{o}(x,R^{\prime})$, then there exists $\xi\in\Phi^{-1}(x)$ such that $z\in H_{o}(\xi,R^{\prime})$. Now if $\zeta\in\Phi^{-1}(x)$ be any other point, then using Result 4.1 we get
$h_{o,\,\zeta}(z)\leq 2\delta+h_{o,\,\xi}(z)<2\delta+(1/2)\log R^{\prime}\leq(1/2)\log R.$
Therefore, for any $\zeta\in\Phi^{-1}(x)$, we have $z\in H_{p}(\zeta,R)$ whence $H^{b}_{o}(x,R^{\prime})\subset H^{s}_{o}(x,R)$. This establishes the claim.
The claim above together with Result 4.2 implies that $\overline{H^{s}_{o}(x,R)}^{G}\cap\partial_{G}\Omega=\{x\}$. For big horospheres, one could change the role of $R,R^{\prime}$ in the claim above, and then the similar argument works. ∎
@@ 5. Continuous extension of biholomorphisms
In this section, we present the proof of our result regarding the continuous extension of biholomorphisms. As an upshot, we shall characterize those bounded simply-connected domains in $\mathbb{C}$ that have property (1.1) in problem ()(*). We first recall a Julia lemma essentially due to Abate.
Result 5.1.
Let $D_{1}\subset\mathbb{C}^{n}$ and $D_{2}\subset\mathbb{C}^{m}$ be bounded domains, and let $f:D_{1}\to D_{2}$ be a holomorphic map. Suppose there exist points $p\in D_{1}$, $q\in D_{2}$, a constant $\alpha>0$, and a sequence $(w_{\nu})\subset D_{1}$ converging to $x\in\partial D_{1}$ such that $f(w_{\nu})\to y\in\partial D_{2}$ and
$\lim_{\nu\to\infty}\!\big(\mathsf{k}_{D_{1}}(p,w_{\nu})-\mathsf{k}_{D_{2}}(q,f(w_{\nu}))\big)\leq\tfrac{1}{2}\log\alpha.$
Then, for every $R>0$,
$f\big(H^{s}_{p}(x,R)\big)\subset H^{b}_{q}(y,\alpha R).$
Abate proved the above result when $D_{1}=D_{2}$ and $p=q$ (see, e.g., [2, Proposition 2.4.15]). The general case follows by the same argument.
@@
Let $D$, $\Omega$ be bounded domains that are complete with respect to the Kobayashi distance. Suppose there exist $o\in D$, $p\in\Omega$ such that
  • for any two points $y_{1}\neq y_{2}\in\partial D$, there exists an $R>0$ such that $H^{b}_{o}(y_{1},R)\cap H^{b}_{o}(y_{2},R)=\emptyset$, and
  • for any $x\in\partial\Omega$ and any $R>0$, $H^{s}_{p}(x,R)\neq\emptyset$.
Then any biholomorphism $\Phi:\Omega\to D$ extends as a continuous map from $\overline{\Omega}$ onto $\overline{D}$.
@@
Let $o\in D$ be as in the statement of the lemma. Observe that — owing to the last part of Result 2.1 — we may assume that $p\in\Omega$ satisfies $\Phi(p)=o$. We claim that, given $x\in\partial\Omega$, let $(w_{n})\subset\Omega$ be a sequence that converges to $x$ such that $\Phi(w_{n})\to y\in\partial D$. Then
$\Phi(H^{s}_{p}(x,R))\subset H^{b}_{o}(y,R)\ \ \ \forall R>0.$ (5.1)
To see this, we note that — since $\Phi$ is an isometry — for each $n$, we have
$\mathsf{k}_{\Omega}(p,w_{n})-\mathsf{k}_{D}(o,\Phi(w_{n}))=\mathsf{k}_{\Omega}(p,w_{n})-\mathsf{k}_{D}(\Phi(p),\Phi(w_{n}))=\mathsf{k}_{\Omega}(p,w_{n})-\mathsf{k}_{D}(p,w_{n})=0.$
Therefore, an application of Result 5.1 with $\alpha=1$, establishes (5.1).
Fix a point $x\in\partial\Omega$, we shall show that the limit
@@
exists. Suppose, to get a contradiction, that the above limit does not exist. This implies that — since $\Phi$ is a biholomorphism and $D$ is bounded — there exist sequences $(z_{n})$ and $(w_{n})$ in $\Omega$ that converge to $x$ and points $y_{1}\neq y_{2}\in\partial D$ such that $\Phi(z_{n})\to y_{1}$, $\Phi(w_{n})\to y_{2}$. Applying (5.1), we get
$\Phi(H^{s}_{o}(x,R))\subset H^{b}_{o}(y_{1},R)\cap H^{b}_{o}(y_{2},R)\ \ \ \forall R>0.$
Since $y_{1}$ and $y_{2}$ are distinct, there exists an $R_{0}>0$ such that $H^{b}_{o}(y_{1},R)\cap H^{b}_{o}(y_{2},R)=\emptyset$ for all $R<R_{0}$. This implies that $H^{s}_{o}(x,R)=\emptyset$ for all $R<R_{0}$ which is a contradiction to our hypothesis whence the result.
It is now routine to check that $\Phi$ extends continuously on $\overline{\Omega}$. ∎
@@
Since small horospheres and big horospheres in $D$ are identical, and $\Psi$ an isometry, an application of Result 5.1 — as in the proof of Lemma 5.2 — implies that given $\zeta\in\partial D$, and $(\zeta_{n})\subset D$, a sequence converging to $\zeta$, such that $\Psi(\zeta_{n})\to y$ for some $y\in\partial\Omega$, we have
@@
Suppose, to get a contradiction, that the above is not true. This implies that — since $\Psi$ is a biholomorphism and $\Omega$ is bounded — there exist sequences $(z_{n}),(w_{n})\subset D$ converging to $\xi$ and $x_{1}\neq x_{2}\in\partial\Omega$ such that $\Psi(z_{n})\to x_{1}$ and $\Psi(w_{n})\to x_{2}$. Applying 5.3, we get